OpenStudy (anonymous):

Factor completely over the real numbers m2(3n + 1)2 − 4m2

OpenStudy (kira_yamato):

m^2 [(3n + 1)^2 - 4] = m^2 [9n^2 + 6n -3] = 3m^2 (3n^2 + 2n - 1) = 3m^2 (3n - 1) (n + 1)